2(x - 3) - 3(x + 5)= 3(x - 2)- 7
Answer by checkley75(3666)

2(x-3)-3(x+5)=3(x-2)-7
2x-6-3x-15=3x-6-7
2x-3x-3x=-6-7+6+15
-6x=8
x=8/-6
x=-4/3 answer.